OpenStudy (anonymous):

Evaluate 3 – 5 · 2 + 8 ÷ 4 – 2 using the correct order of operations. -1 -3 -7 -10

OpenStudy (solomonzelman):

Multiplication and division first then addition and subtraction.

OpenStudy (anonymous):

5x2=10. 8/4=2. 3-10=-7. 2-2=0. -7+0=-7

OpenStudy (anonymous):

-7 @SolomonZelman

OpenStudy (anonymous):

:P I got the answer

OpenStudy (solomonzelman):

\[3-10+4-2\]\[-7+4-2\]\[-3-2\]\[-5\]

OpenStudy (solomonzelman):

And where is -7 from?

OpenStudy (anonymous):

you do realize that -5 is not an option right @SolomonZelman

OpenStudy (anonymous):

Plus, if you plug it into a calculator, you will get -7

OpenStudy (anonymous):

I did what you said. There also isn't a -5 option :(

OpenStudy (anonymous):

Do I get a medal? I got it right ;)

OpenStudy (anonymous):

no you got 2 medals which means double right lol

OpenStudy (solomonzelman):

I did make a mistake, this is really silly, sorry.

OpenStudy (anonymous):

I love you guys!

OpenStudy (solomonzelman):

it's +2 not +4, so then -7

OpenStudy (anonymous):

@wangkan07 lol :)

OpenStudy (anonymous):

*pat on the back* we love you too @wangkan07

OpenStudy (anonymous):

@shelbygt520 lol ;)

OpenStudy (anonymous):

well that's new an lol? thats rare

OpenStudy (anonymous):

:P